• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 14:12:33 on October 7,2002

    Entry number 87562

    keyword=cryotarget cool-down completed

    At about 1 pm, the cryotarget group (Dave) completed the cool-down of
    the LD2 target (in loop 2 now). Temperature is 22K, pressure is about 23 psi.
    They swaped LH2 with LD2 by purge with Helium gas, so that it was done
    without fully warm up everything.